QUESTION: If two(2) men can build two(2) engines in two(2) hours, then how many men will it take to build eighteen(18) engines in six(6) hours?
I do not know how to set up this problem to be solved. Thank you for your services
Sincerely,
Amerikaner

Answer by ankor@dixie-net.com(22740)

QUESTION: If two(2) men can build two(2) engines in two(2) hours, then how many men will it take to build eighteen(18) engines in six(6) hours?
:
One way to do this, is by finding the the man-hours required to build one engine>
:
It says:
2*2 = 4 man-hours to build two engines, therefore:
it takes 2 man-hours to build one engine
:
Let x = no. of men required to build 18 engines in 6 hrs
:
Man-hrs = no of engines * 2 man-hrs/per engine
6x = 18*2
6x = 36
x = 36/6
x = 6 men to build 18 engines in 6 hrs
:
Did this make sense to you?